OpenStudy (anonymous):

The demand curve for original Iguanawoman comics is given by q=(343-p)^2/100 (0

OpenStudy (dumbcow):

haha an econ question. ok price elasticity is looking at the slope of the demand curve to see how responsive the price is to a change in quantity. E = price/quantity * q'(p) p = 33 then q = 961 derivative of function using chain rule -> q'(p) = -(343-p)/50 -> E = (33/961)*-(343-33)/50 = -0.2129

OpenStudy (dumbcow):

for part b) revenue is price * quantity -> R(p) = p*(343-p)^2/100 to max revenue differentiate using product rule and set equal to 0 R'(p) = (343-p)^2/100 + -p(343-p)/50 = 0 solve for p -> p = 343/3 For part c) plug p=343/3 back into revenue function

OpenStudy (anonymous):

didn't work :/